OpenStudy (anonymous):

$1,600 principal earning 7%, compounded semi-annually, after 33 years. solve and explain

OpenStudy (campbell_st):

the number of time periods needs to double... from 33 years to 66 half years... to match the compounding period. the formula is \[A = P(1 + \frac{r}{100})^n\] P = principal, r = interest rate and n = number of time periods the only pert I am unclear on is the interest rate.... is it 7% per annum or 7% per half year? the calculations are HALF YEAR at 7% \[A = 1600(1+\frac{7}{100})^{66}\] HALF YEAR at 3.5% you need to change the annual interest to half yearly 7/2 = 3.5 \[A = 1600(1 + \frac{3.5}{100})^{66}\] just calculate you answer